OpenStudy (anonymous):

he coefficient of the product of (-5xy2) and (-4x2y) is . The exponent of x in the product is , and the exponent of y is .

OpenStudy (solomonzelman):

at first, you would just need to find the product and then to label everything in it.

OpenStudy (solomonzelman):

The question is, \(\large\color{black}{ (-5xy^2)\times(-4x^2y) }\). Just answer a couple of questions to yourself./ What is \(\large\color{black}{ (-5)\times(-4) }\) ? What is \(\large\color{black}{ x\times x^2 }\) ? What is \(\large\color{black}{ y^2 \times y }\) ?

OpenStudy (solomonzelman):

then just put it all together. After you find the product: The exponent of x, is the power to which x will be raised. The exponent of y, is the power to which y will be raised. The coefficient, is the constant/number times which the variable or variables, is/are multiplied by.

OpenStudy (solomonzelman):

Hope this helped....